The measure of a supplement of an angle is four more than three times the angle. How much is the angle and the supplement?
I got: 44 degrees, 136 degrees
Not sure if this is correct.
Thanks~

Answer by tutorcecilia(2152)

Correct!
Angle x + Angle y =180
Angle y = 3x+4, so
x+(3x+4)=180 [solve for x]
3x+4=180
3x=180-4
3x=176
x=44
.
Plugging in the values
44+[3(44)+4]=180
44+(132+4)=180
180=180 [checks out]